Two vectors of equal magnitude "P" are inclined at some angle such that the difference in the magnitude of resultant and magnitude of either of the vectors is 0.732 times either of the vectors. If the angle between them is increased to half of its initial 
Find the value of difference of the vectors

Grade:

2 Answers

Mallikarjun Maram
9 years ago
Dear Student, The problem statement is not clear.  How can you increase the angle to half of its original value?  Please verify and repost if needed.
anudeep mahesh
6 years ago
Given R-0.732p=p , |a|=p,|b|=p R=1.732p R^2= 3p^2 2p^2(1+cosΦ)=3p^2 CosΦ=1/2 Φ=60°Now, new resulatant makes angle (Φ+1/2Φ) then 60+30=90° Ř=√2p^2(1-cos90°) Ř=√2p Ans:√2p